OpenStudy (anonymous):

Solve: 2 sin x - 1 = 0 for 0

OpenStudy (anonymous):

sinx = 1/2 => x = 30° or 150° sin2x/cos2x = 1 => tan2x = 1 => 2x = 45°, 180° + 45°, 360° + 45° and 540° + 45°..............45 + 180n, for n = 0, 1, 2,..... i.e. x = 22.5°, 112.5°, 202.5°, 292.5°

OpenStudy (anonymous):

By manipulating the equation, you get: sin x = 1/2 acute angle = sin^-1 (1/2) = 30degrees (I presume the domain you are given is 0 to 360 degrees for x.) Therefore, x = 30 degrees, 180-30degrees(because this angle also appears in the 'sine domain'.) = 30 degrees, 150 degrees